OpenStudy (anonymous):

Consider the line y=5/3x-7 Find the equation of the line that is parallel to this line and passes through the point (3,5) . Find the equation of the line that is perpendicular to this line and passes through the point (3,5)

OpenStudy (nory):

Parallel lines have the same slopes. So the slope of the parallel line would be 5/3. Now, because we have the slope and a point on the line, we can use the point-slope format. Do you remember that one?

OpenStudy (anonymous):

@Nory um y2-y1/x2-x1 right

OpenStudy (nory):

That is one form of it. Can you put the slope and the point into that equation to get the equation?

OpenStudy (anonymous):

I can put the points but I don't know about the slope @Nory

OpenStudy (anonymous):

@Nory wait is it y=5/3+5?

OpenStudy (nory):

Almost. I believe that it would be y = 5/3x. Because we have the point (3,5) and 5 = 5/3 * 3.

OpenStudy (anonymous):

oh okay @Nory so y=5/3x+5 is my answer for the first question?

OpenStudy (nory):

I think it would be y = 5/3x...where did you get the 5 from?

OpenStudy (anonymous):

(3,5) I took the 5 as the y intercept

OpenStudy (nory):

Five is not the y-intercept.

OpenStudy (anonymous):

So then how would I get it? @Nory

OpenStudy (anonymous):

@Nory

OpenStudy (nory):

Do you know the equation y - y1 = m(x - x1) ?

OpenStudy (anonymous):

yes @Nory

OpenStudy (nory):

Well use that. Use (3,5) for your (x1, y1) and 5/3 for your m.

OpenStudy (anonymous):

I get y=5/3x+0 @Nory

OpenStudy (nory):

Good job! And there's your answer.

OpenStudy (anonymous):

I even keep the +0? @Nory

OpenStudy (anonymous):

and also that's for the first question?

OpenStudy (nory):

Now you know how to do the second one, almost. Except that perpendicular lines have opposite reciprocal slopes. So if the original line has a slope of 5/3, the perpendicular has a slope of -3/5.

OpenStudy (nory):

Yes it is. You can keep the 0 or not; most people don't require that you keep the 0 on.

OpenStudy (anonymous):

Okay! Thank you for the help! @Nory

OpenStudy (nory):

No problem. Anytime.